OpenStudy (anonymous):

An aquarium has a volume of 60,480in^3. The aquarium has a height of 36 inches and a length of 4 more inches than twice the width. Find the length and width of the aquarium.

OpenStudy (anonymous):

x= width y= length

jimthompson5910 (jim_thompson5910):

"length of 4 more inches than twice the width" means y = x+4

jimthompson5910 (jim_thompson5910):

Volume V = L*W*H V = y*x*36 V = (x+4)*x*36 V = 36x^2 + 144x 60480 = 36x^2 + 144x keep going to solve for x

OpenStudy (anonymous):

4 more than twice the width, wouldn't that be y = 2x + 4?

jimthompson5910 (jim_thompson5910):

oh right, my bad

jimthompson5910 (jim_thompson5910):

so it would be V = (2x+4)*x*36